# Physics-informed machine learning for combustion: A review

### v1

@@ 2.3.2 Optimization process
Sampling. Sampling corresponds to the modification of $\mathcal{S}$ in Equations (2.5) and (2.6). Basic sampling methods include the uniform, random, and Latin hypercube sampling. To account for the real-time information on the distribution of residuals, thus enhancing the convergence, some adaptive sampling algorithms have been proposed 167, 168, 169.
Loss weights. Since PINNs belong to the multi-task learning, the imbalance/conflict among different loss terms is a common failure mode in PINNs 170. Therefore, the modification of the loss weights, i.e., $w$ in Equation (2.8), is important. Some examples include designing learning rate annealing 170, utilizing the neural tangent kernel (NTK) theory 171, the self-adaptive loss balanced PINNs (lbPINNs) 172, and the self-adaptive PINNs (SA-PINNs) 173. For time-dependent problems, it is important to respect the temporal causality, which can be achieved by assigning different weights to the residual points at different time 174, 175.
Loss definition. Some algorithms add loss terms to Equation (2.8) to enhance regularization. Examples are the gradient-enhanced PINNs (gPINNs) 176 and region optimized PINNs (RoPINNs) 177. Some algorithms replace the $L^{2}$ loss in Equations (2.5) to (2.7) with $L^{\infty}$ loss 178 and $L^{p}$ loss 128.
@@
Domain decomposition. Domain decomposition is important when the solution exhibits distinct behaviours in different domains. For time-dependent problems, the time-marching strategy that decomposes the temporal domain can improve the long-time prediction accuracy. In most cases, separate NNs are assigned to different subdomains and the interface conditions between adjacent subdomains need to be considered. Examples of PINNs using domain decomposition include distributed PINNs (DPINNs) 181, conservative PINNs (cPINNs) 182, extended PINNs (XPINNs) 183, parareal PINNs (PPINNs) 184, parallel PINNs 185, spatiotemporal parallel PINNs (STPINNs) 186, augmented PINNs (APINNs) 187, finite basis PINNs (FBPINNs) 188, 189, and dtPINNs 175.
Optimizer. The most used optimizers in Equations (2.9) and (2.10) are Adam 190 and L-BFGS 191. Some improved optimizers for PINNs include the MultiAdam 192 and the NNCG 193.
@@ 2.3.3 PDE system
Variational form. The variational form of PDEs is essential for traditional FEM solvers and has been applied to PINNs 196, 197. Similar to the variational form, the energy form of PDEs is important in solid mechanics 94 and has also been adopted in PINNs 198, 199, 200, 201. An advantage of variational form is the reduction of derivative order compared to the original PDEs, thus accelerating the computation.
Other formal transformations. For example, the stream function can be used in 2D incompressible flow problems 202, 170. In this way, the continuity/incompressibility equation can be exactly satisfied but higher-order derivatives are introduced. Besides, the common velocity-pressure form of incompressible NS equations can be transformed to the vorticity-velocity form 203. For inverse problems with sparse and noisy data, the filtered PDEs can serve as an effective substitute for the original PDEs 204.
Other DE types. PINNs have also been applied to non-traditional types of DEs, such as fractional DEs 205 and stochastic DEs 154, 155.
@@ 2.3.4 Other important topics
Hard constraint. The classical PINNs impose the physical constraints in a soft manner, so there is still a possibility that physical constraints will be violated if the corresponding residuals do not converge. In contrast, hard constraints can ensure the physical laws are exactly satisfied, thus reducing the number of loss terms in most cases. Hard constraints are also referred to as physics "-embedded", "-encoded", "-enforced", "-preserved" in many papers. For IC/BCs, they can be enforced by input and output transformations as mentioned in Section 2.3.1. BCs can also be strictly encoded in discrete PINNs via padding operations 99, 121. For PDEs, one example is to enforce the continuity equation using the stream function as mentioned in Section 2.3.3. PDEs can also be constrained by linear projection 111. For discrete PINNs, PDEs can be encoded using convolutional kernels, such as the physics-encoded RCNN (PeRCNN) 73 and the PDE-preserved NN (PPNN) 74, and can be constrained using traditional numerical methods/solvers 114, 115, 113, 116. In a broad sense, any explicitly written physical laws that are always satisfied can be regarded as hard constraints. For combustion, the species and element conservations can also be enforced. See Section 4.2 and Equation (5.2) for details. The laws of chemical kinetics can also be embedded, which will be introduced in Section 4.3. The choice between soft and hard constraints will be discussed in Section 8.2.
@@
Uncertainty quantification (UQ). UQ is important for reliable modeling of practical systems due to the intrinsic incompleteness of data and model as well as the discrepancies between them 213. UQ has been applied in PINNs for quantifying both data uncertainty 214, 215, 216 and model uncertainty 217.
Theory. The total error of PINNs can be decomposed into three parts 218: approximation error, estimation error, and optimization error. Many studies have estimated the errors of PINNs, such as for forward problems 219, inverse problems 220, Kolmogorov PDEs 221, and NS equations 222. The aforementioned NTK theory for PINNs 171 and ill-conditioning in PINNs are also theoretical advances in PIML.
@@ 2.4 Related applications
PIML has been widely applied to fluid mechanics, ranging from incompressible flows 202, 203 to supersonic flows 223, 224, 225, from laminar flows 202 to turbulent flows 226, 227, 228, 229, 230 and transitional flows 231, from single-phase flows to two-phase flows 232, 233, 234, from internal flows to external flows such as boundary layer flows 235, 236, 85, 237, and from analytical and computational fluid mechanics 203 to experimental fluid mechanics 238, 239, 240, 241, 242, 243, 204. Compared to forward problems, solving inverse problems using PINNs serves as a new paradigm in fluid mechanics which compensates for the shortcomings of traditional algorithms. One representative method is the hidden fluid mechanics (HFM) 244, which can reconstruct the velocity and pressure fields based on the observations of concentration. Flow field reconstruction is quite important for experimental fluid mechanics because the data are always sparse, noisy, and partial. For that, PINNs have been widely applied to enhance the experimental data obtained from particle image velocimetry (PIV) 239, 240, particle tracking velocimetry (PTV) 239, 243, and other measurements 238, 241, 242, 204. Inverse problems are also important for turbulence modeling based on Reynolds-averaged Navier-Stokes (RANS) equations and large eddy simulations (LES), where PINNs have been applied to both the task of field reconstruction 226, 227, 228 and parameter inference 229, 230. The parametric/surrogate modeling mentioned in Section 2.3.4 has also been applied to fluid mechanics, including the cardiovascular flows 245, Euler equations 225, and flows around airfoils 246, 247, 248. For more details of PINNs for fluid mechanics, readers can refer to these reviews 87, 88, 89. In addition to pure fluid flows, PINNs have also been applied to heat transfer 249, 250, 251, mass transfer, and multi-species flows 252. These are all subprocesses of combustion reacting flows. For a detailed introduction, please refer to this review 91.
@@ 3.5 Fluid dynamics
Equations (3.38) to (3.42) are the unsimplified governing equations for reacting flows and are used for direct numerical simulation (DNS). To save computational resources, simplified models are used in many cases. For example, the concepts of reaction progress variable ($c$) and mixture fraction ($Z$) can be derived in the contexts of premixed and non-premixed combustion, respectively 5. Under certain assumptions, their governing equations can also be derived, which are fewer than the original equations. For turbulent combustion, the governing equations for Reynolds-averaged Navier-Stokes simulation (RANS) and large eddy simulation (LES) are also different from Equations (3.38) to (3.42). For two types of turbulent combustion models, namely the transport probability density function (TPDF) method 253 and the flamelet-like method 254, 13, including the flamelet generated manifolds (FGM) 255, 15 and flamelet/progress variable (FPV) method 256, 257, the corresponding governing equations are also different from Equations (3.38) to (3.42). For the sake of brevity, these simplified methods are not described in detail here. For details, readers can refer to the corresponding references.
@@ 4 PIML for combustion chemical kinetics (0D combustion)
Ji 258, 2021.8 stiff-PINN Solve stiff ODEs Loss of ODEs URL
De Florio 259, 2022.6 Solve stiff ODEs Loss of ODEs
Wang 260, 2022.7 Solve stiff ODEs Loss of ODEs
Galaris 261, 262, 2021.8 PI-RPNN Solve stiff ODEs Loss of ODEs URL
Li 263, 2021.10 Solve ODEs Loss of ODEs
Weng 264, 2022.11 M-PINN Solve stiff ODEs (inverse) Loss of ODEs
Gusmao 265, 2022.4 KINN Solve ODEs (forward and inverse) Loss of ODEs, $\mathcal{C}_{\mathrm{spe}}$ by trigonometric transformation URL
Bibeau 266, 2023.12 Solve inverse ODEs (parameter inference) Loss of ODEs URL
Almeldein 267, 2023.7 Surrogate model Loss of $\mathcal{C}_{\mathrm{ele}}$
Zhang 268, 2024.8 CRK-PINN Surrogate model Losses of ODEs, $\mathcal{C}_{\mathrm{ele}}$, $\mathcal{C}_{\mathrm{spe}}$, and $\mathcal{C}_{\mathrm{ene}}$
Wan 51, 2020.7 Surrogate model of reduced chemistry $\mathcal{C}_{\mathrm{ele}}$ by correction
Wang 113, 2025.3 ANN-hard Surrogate model $\mathcal{C}_{\mathrm{ele}}$ by correction, $\mathcal{C}_{\mathrm{ene}}$ by using equations Dataset
Kumar 102, 103, 2023.12 Surrogate model Losses of $\mathcal{C}_{\mathrm{ele}}$ and $\mathcal{C}_{\mathrm{spe}}$
Zanardi 104, 2021.12 Surrogate model Loss of ODEs
Zanardi 105, 2023.9 CG-DeepONet Surrogate model Loss of ODEs, Boltzmann layer, hierarchical architecture
Weng 269, 2024.11 EFNO Surrogate model Losses of $\mathcal{C}_{\mathrm{ele}}$ and $\mathcal{C}_{\mathrm{spe}}$
Readshaw or Ding 270, 271, 272, 2021.3 HFRD-MMP Surrogate model $\mathcal{C}_{\mathrm{spe}}$ by normalization
Readshaw 273, 2023.4 Surrogate model $\mathcal{C}_{\mathrm{ele}}$ by recalculating residual species
Salunkhe 274, 275, 2022.6 ChemTab Surrogate model (PVG and MA) Constraints on the projection matrix and progress variables
Vijayarangan 276, 2023.11 Surrogate model by NODE Dynamics-informed discovery of latent space
Kumar 277, 278, 2023.11 PC-NODE Surrogate model by NODE Loss of $\mathcal{C}_{\mathrm{ele}}$
Kumar 279, 2024.12 Phy-ChemNODE Surrogate model by NODE Dynamics-informed discovery of latent space, loss of $\mathcal{C}_{\mathrm{ele}}$
Koenig 280, 2025.7 ChemKAN Surrogate model by NODE Architecture that mimics the chemical kinetics, loss of $\mathcal{C}_{\mathrm{ele}}$
Sorourifar 117, 2023.9 r-NODE, k-NODE Surrogate model by NODE Embedded stoichiometric and rate information
Fedorov 118, 2023.10 KC-NODE Surrogate model by NODE Embedded stoichiometry and rate expression URL
Kircher 119, 120, 2024.2 GRNN Surrogate model by NODE Embedded stoichiometry and thermodynamics
Ji 75, 2021.1 CRNN Discover kinetics Embedded rate expression URL
Su 281, 2023.3 Learn kinetics parameters Embedded rate expression
Doeppel 282, 2024.7 AC-CRNN Discover kinetics Embedded rate expression, $\mathcal{C}_{\mathrm{ele}}$ by a NN layer
@@ 4.1 Solving ODEs of chemical kinetics
where $\lambda_{j}\in\mathbb{C}$ are the eigenvalues of $\mathbf{J}$. The reciprocals of the eigenvalues can represent the time scales, $\tau$, of the system 276. To validate the algorithms (traditional and PINNs) for solving stiff ODEs, some classical problems are widely tested, including the ROBER and POLLU problems, which consist of 3 species and 3 reactions, and 20 species and 25 reactions, respectively. As an example, the ROBER problem describes the following reaction:
@@
Ji et al. 258 first proposed the stiff-PINNs (shown in Figure 5), using PINNs to solve stiff chemical kinetic ODEs, where the quasi-steady-state assumption (QSSA) was employed to reduce the stiffness of the system. In QSSA, QSS species are identified by assuming that their net production rates are negligible compared to their consumption and production rates, then the ODEs of QSS species become differential-algebraic equations (DAEs). For example, $B$ can be chosen as the QSS species in the ROBER problem, so its algebraic equation can be obtained by assuming $\mathrm{d}y_{2}/\mathrm{d}t=0$. Because QSS species are usually radicals and unstable intermediates with short time scales, the elimination of their ODEs can greatly reduce the stiffness. The stiff-PINNs were used to solve the ROBER and POLLU problems with one and ten QSS species, respectively, and yielded satisfactory results. De Florio et al. 259 used X-TFC 133 (see Section 2.3.2) to solve stiff chemical kinetic ODEs, where the TFC can guarantee the ICs are exactly satisfied, as illustrated in Figure 6. The temporal domain was decomposed into several intervals to deal with the long-time issue. The X-TFC showed great performance for 4 tested problems, and even outperformed the traditional numerical methods in some cases. Wang et al. 260 used PINNs to solve the ROBER problem. Hard constraints on ICs via logarithmic functions were employed, yielding accurate results. Galaris et al. 261 proposed the physics-informed random projection NNs (PI-RPNNs) to solve stiff ODEs, where the single-layer NN had fixed weights between the input and hidden layer so only the parameters from the hidden to the output layer needed to be optimized. The loss functions were minimized by the Gauss-Newton method and the hard constraints on ICs were designed. Four stiff ODE problems were tested and PI-RPNN yielded good accuracy and can outperform the traditional solvers in the cases with steep gradients. Later, Fabiani et al. 262 extended the PI-RPNNs to other stiff ODEs and three index-1 DAEs. For more practical applications, Li et al. 263 applied PINNs to solve the coal gasification chemical kinetic problems. The unreacted-core shrinking model was adopted for the gas-solid reactions, the temperature was simplified to be constant, and the ICs were hardcoded. Results showed that PINNs could accurately predict the production rates of the gaseous species.
@@
Figure 5: Schematic of the stiff-PINN for the ROBER problem. Reprinted from Reference 258, with permission.
@@
Figure 6: Schematic of the X-TFC algorithm to solve stiff chemical kinetics equations. Reprinted from Reference 259, with permission.
@@
Although the above studied focus on solving the chemical kinetic ODEs in a forward way without data supervision, inverse problems can also be solved when the available data are utilized. Weng et al. 264 developed the multiscale PINNs for stiff chemical kinetic problems, where the species with different time scales were classified and trained by multiple NNs. By introducing ground truth data points, the proposed method showed superior ability on solving three stiff chemical kinetic problems. Gusmao et al. 265 proposed the kinetics-informed NNs (KINNs) to solve forward and inverse problems of chemical kinetics, as shown in Figure 7. Practical species and reaction types were considered, including three types of species (gas, absorbed molecules, and surface radicals) and five types of reactions (denoted by g, a, d, c, and s). ICs were enforced by an IC operator coupled with a shallow NN and species conservation was enforced by trigonometric transformations. In forward problems, no data were provided, while in inverse problems, synthetic data were available and three scenarios were designed based on the information about chemical intermediates and adsorbed molecules, and the presence of noise. The unknown kinetic parameters were optimized together with the NN parameters in inverse problems. The results showed that KINNs can accurately solve the ODEs in forward problems and estimate the kinetic parameters in inverse problems. Later, KINNs were applied to model the data from temporal analysis of products 283 and extended to a robust version by incorporating the maximum likelihood approach 284. Bibeau et al. 266 used PINNs to predict the kinetics of biodiesel production in microwave reactors. The system contained five species and three reactions. The Arrhenius equation was linearized and the temperature evolution was simplified to a first-order ODE, resulting in 15 unknown parameters related to the reaction kinetics and energy balance, which were estimated by PINNs from experimental data. Results showed that the reaction kinetics of the biodiesel process could be identified successfully by PINNs with satisfactory extrapolation ability.
@@
Figure 7: Kinetics-informed neural networks. Reprinted from Reference 265, with permission.

### v2

@@ 2.3.2 Optimization process
Sampling. Sampling corresponds to the modification of $\mathcal{S}$ in Equations (2.5) and (2.6). Basic sampling methods include the uniform, random, and Latin hypercube sampling. To account for the real-time information on the distribution of residuals, thus enhancing the convergence, some adaptive sampling algorithms have been proposed 170, 171, 172.
Loss weights. Since PINNs belong to the multi-task learning, the imbalance/conflict among different loss terms is a common failure mode in PINNs 173. Therefore, the modification of the loss weights, i.e., $w$ in Equation (2.8), is important. Some examples include designing learning rate annealing 173, utilizing the neural tangent kernel (NTK) theory 174, the self-adaptive loss balanced PINNs (lbPINNs) 175, and the self-adaptive PINNs (SA-PINNs) 176. For time-dependent problems, it is important to respect the temporal causality, which can be achieved by assigning different weights to the residual points at different time 177, 178.
Loss definition. Some algorithms add loss terms to Equation (2.8) to enhance regularization. Examples are the gradient-enhanced PINNs (gPINNs) 179 and region optimized PINNs (RoPINNs) 180. Some algorithms replace the $L^{2}$ loss in Equations (2.5) to (2.7) with $L^{\infty}$ loss 181 and $L^{p}$ loss 131.
@@
Domain decomposition. Domain decomposition is important when the solution exhibits distinct behaviours in different domains. For time-dependent problems, the time-marching strategy that decomposes the temporal domain can improve the long-time prediction accuracy. In most cases, separate NNs are assigned to different subdomains and the interface conditions between adjacent subdomains need to be considered. Examples of PINNs using domain decomposition include distributed PINNs (DPINNs) 184, conservative PINNs (cPINNs) 185, extended PINNs (XPINNs) 186, parareal PINNs (PPINNs) 187, parallel PINNs 188, spatiotemporal parallel PINNs (STPINNs) 189, augmented PINNs (APINNs) 190, finite basis PINNs (FBPINNs) 191, 192, and dtPINNs 178.
Optimizer. The most used optimizers in Equations (2.9) and (2.10) are Adam 193 and L-BFGS 194. Some improved optimizers for PINNs include the MultiAdam 195 and the NNCG 196.
@@ 2.3.3 PDE system
Variational form. The variational form of PDEs is essential for traditional FEM solvers and has been applied to PINNs 199, 200. Similar to the variational form, the energy form of PDEs is important in solid mechanics 97 and has also been adopted in PINNs 201, 202, 203, 204. An advantage of variational form is the reduction of derivative order compared to the original PDEs, thus accelerating the computation.
Other formal transformations. For example, the stream function can be used in 2D incompressible flow problems 205, 173. In this way, the continuity/incompressibility equation can be exactly satisfied, but higher-order derivatives are introduced. Besides, the common velocity-pressure form of incompressible NS equations can be transformed to the vorticity-velocity form 206. For inverse problems with sparse and noisy data, the filtered PDEs can serve as an effective substitute for the original PDEs 207.
Other DE types. PINNs have also been applied to non-traditional types of DEs, such as fractional DEs 208 and stochastic DEs 157, 158.
@@ 2.3.4 Other important topics
Hard constraint. The classical PINNs impose the physical constraints in a soft manner, so there is still a possibility that physical constraints will be violated if the corresponding residuals do not converge. In contrast, hard constraints can ensure the physical laws are exactly satisfied, thus reducing the number of loss terms in most cases. Hard constraints are also referred to as physics "-embedded", "-encoded", "-enforced", "-preserved" in many papers. For IC/BCs, they can be enforced by input and output transformations as mentioned in Section 2.3.1. BCs can also be strictly encoded in discrete PINNs via padding operations 102, 124. For PDEs, one example is to enforce the continuity equation using the stream function as mentioned in Section 2.3.3. PDEs can also be constrained by linear projection 114. For discrete PINNs, PDEs can be encoded using convolutional kernels, such as the physics-encoded RCNN (PeRCNN) 76 and the PDE-preserved NN (PPNN) 77, and can be constrained using traditional numerical methods/solvers 117, 118, 116, 119. In a broad sense, any explicitly written physical laws that are always satisfied can be regarded as hard constraints. For combustion, the species and element conservations can also be enforced. See Section 4.2 and Equation (5.2) for details. The laws of chemical kinetics can also be embedded, which will be introduced in Section 4.3. The choice between soft and hard constraints will be discussed in Section 8.2.
@@
Uncertainty quantification (UQ). UQ is important for reliable modeling of practical systems due to the intrinsic incompleteness of data and model as well as the discrepancies between them 216. UQ has been applied in PINNs for quantifying both data uncertainty 217, 218, 219 and model uncertainty 220.
Theory. The total error of PINNs can be decomposed into three parts 221: approximation error, estimation error, and optimization error. Many studies have estimated the errors of PINNs, such as for forward problems 222, inverse problems 223, Kolmogorov PDEs 224, and NS equations 225. The aforementioned NTK theory for PINNs 174 and ill-conditioning in PINNs are also theoretical advances in PIML.
@@ 2.4 Related applications
PIML has been widely applied to fluid mechanics, ranging from incompressible flows 205, 206 to supersonic flows 226, 227, 228, from laminar flows 205 to turbulent flows 229, 230, 231, 232, 233 and transitional flows 234, from single-phase flows to two-phase flows 235, 236, 237, from internal flows to external flows such as boundary layer flows 238, 239, 88, 240, and from analytical and computational fluid mechanics 206 to experimental fluid mechanics 241, 242, 243, 244, 245, 246, 207. Compared to forward problems, solving inverse problems using PINNs serves as a new paradigm in fluid mechanics which compensates for the shortcomings of traditional algorithms. One representative method is the hidden fluid mechanics (HFM) 247, which can reconstruct the velocity and pressure fields based on the observations of concentration. Flow field reconstruction is quite important for experimental fluid mechanics because the data are always sparse, noisy, and partial. For that, PINNs have been widely applied to enhance the experimental data obtained from particle image velocimetry (PIV) 242, 243, particle tracking velocimetry (PTV) 242, 246, and other measurements 241, 244, 245, 207. Inverse problems are also important for turbulence modeling based on Reynolds-averaged Navier-Stokes (RANS) equations and large eddy simulations (LES), where PINNs have been applied to both the task of field reconstruction 229, 230, 231 and parameter inference 232, 233. The parametric/surrogate modeling mentioned in Section 2.3.4 has also been applied to fluid mechanics, including the cardiovascular flows 248, Euler equations 228, and flows around airfoils 249, 250, 251. For more details of PINNs for fluid mechanics, readers can refer to these reviews 90, 91, 92. In addition to pure fluid flows, PINNs have also been applied to heat transfer 252, 253, 254, mass transfer, and multi-species flows 255. These are all subprocesses of combustion reacting flows. For a detailed introduction, please refer to this review 94.
@@ 3.5 Fluid dynamics
Equations (3.38) to (3.42) are the unsimplified governing equations for reacting flows and are used for direct numerical simulation (DNS). To save computational resources, simplified models are used in many cases. For example, the concepts of reaction progress variable ($c$) and mixture fraction ($Z$) can be derived in the contexts of premixed and non-premixed combustion, respectively 5. Under certain assumptions, their governing equations can also be derived, which are fewer than the original equations. For turbulent combustion, the governing equations for Reynolds-averaged Navier-Stokes simulation (RANS) and large eddy simulation (LES) are also different from Equations (3.38) to (3.42). For two types of turbulent combustion models, namely the transport probability density function (TPDF) method 256 and the flamelet-like method 257, 13, including the flamelet generated manifolds (FGM) 258, 15 and flamelet/progress variable (FPV) method 259, 260, the corresponding governing equations are also different from Equations (3.38) to (3.42). For the sake of brevity, these simplified methods are not described in detail here. For details, readers can refer to the corresponding references.
@@ 4 PIML for combustion chemical kinetics (0D combustion)
Ji 261, 2021.8 stiff-PINN Solve stiff ODEs Loss of ODEs URL
De Florio 262, 2022.6 Solve stiff ODEs Loss of ODEs
Wang 263, 2022.7 Solve stiff ODEs Loss of ODEs
Galaris 264, 265, 2021.8 PI-RPNN Solve stiff ODEs Loss of ODEs URL
Li 266, 2021.10 Solve ODEs Loss of ODEs
Weng 267, 2022.11 M-PINN Solve stiff ODEs (inverse) Loss of ODEs
Gusmao 268, 2022.4 KINN Solve ODEs (forward and inverse) Loss of ODEs, $\mathcal{C}_{\mathrm{spe}}$ by trigonometric transformation URL
Bibeau 269, 2023.12 Solve inverse ODEs (parameter inference) Loss of ODEs URL
Almeldein 270, 2023.7 Surrogate model Loss of $\mathcal{C}_{\mathrm{ele}}$
Zhang 271, 2024.8 CRK-PINN Surrogate model Losses of ODEs, $\mathcal{C}_{\mathrm{ele}}$, $\mathcal{C}_{\mathrm{spe}}$, and $\mathcal{C}_{\mathrm{ene}}$
Wan 54, 2020.7 Surrogate model of reduced chemistry $\mathcal{C}_{\mathrm{ele}}$ by correction
Wang 116, 2025.3 ANN-hard Surrogate model $\mathcal{C}_{\mathrm{ele}}$ by correction, $\mathcal{C}_{\mathrm{ene}}$ by using equations Dataset
Kumar 105, 106, 2023.12 Surrogate model Losses of $\mathcal{C}_{\mathrm{ele}}$ and $\mathcal{C}_{\mathrm{spe}}$
Zanardi 107, 2021.12 Surrogate model Loss of ODEs
Zanardi 108, 2023.9 CG-DeepONet Surrogate model Loss of ODEs, Boltzmann layer, hierarchical architecture
Weng 272, 2024.11 EFNO Surrogate model Losses of $\mathcal{C}_{\mathrm{ele}}$ and $\mathcal{C}_{\mathrm{spe}}$
Readshaw or Ding 273, 274, 275, 2021.3 HFRD-MMP Surrogate model $\mathcal{C}_{\mathrm{spe}}$ by normalization
Readshaw 276, 2023.4 Surrogate model $\mathcal{C}_{\mathrm{ele}}$ by recalculating residual species
Salunkhe 277, 278, 2022.6 ChemTab Surrogate model (PVG and MA) Constraints on the projection matrix and progress variables
Vijayarangan 279, 2023.11 Surrogate model by NODE Dynamics-informed discovery of latent space
Kumar 280, 281, 2023.11 PC-NODE Surrogate model by NODE Loss of $\mathcal{C}_{\mathrm{ele}}$
Kumar 282, 2024.12 Phy-ChemNODE Surrogate model by NODE Dynamics-informed discovery of latent space, loss of $\mathcal{C}_{\mathrm{ele}}$
Koenig 283, 2025.7 ChemKAN Surrogate model by NODE Architecture that mimics the chemical kinetics, loss of $\mathcal{C}_{\mathrm{ele}}$
Sorourifar 120, 2023.9 r-NODE, k-NODE Surrogate model by NODE Embedded stoichiometric and rate information
Fedorov 121, 2023.10 KC-NODE Surrogate model by NODE Embedded stoichiometry and rate expression URL
Kircher 122, 123, 2024.2 GRNN Surrogate model by NODE Embedded stoichiometry and thermodynamics
Ji 78, 2021.1 CRNN Discover kinetics Embedded rate expression URL
Su 284, 2023.3 Learn kinetics parameters Embedded rate expression
Doeppel 285, 2024.7 AC-CRNN Discover kinetics Embedded rate expression, $\mathcal{C}_{\mathrm{ele}}$ by a NN layer
@@ 4.1 Solving ODEs of chemical kinetics
where $\lambda_{j}\in\mathbb{C}$ are the eigenvalues of $\mathbf{J}$. The reciprocals of the eigenvalues can represent the time scales, $\tau$, of the system 279. To validate the algorithms (traditional and PINNs) for solving stiff ODEs, some classical problems are widely tested, including the ROBER and POLLU problems, which consist of 3 species and 3 reactions, and 20 species and 25 reactions, respectively. As an example, the ROBER problem describes the following reaction:
@@
Ji et al. 261 first proposed the stiff-PINNs (shown in Figure 5), using PINNs to solve stiff chemical kinetic ODEs, where the quasi-steady-state assumption (QSSA) was employed to reduce the stiffness of the system. In QSSA, QSS species are identified by assuming that their net production rates are negligible compared to their consumption and production rates, then the ODEs of QSS species become differential-algebraic equations (DAEs). For example, $B$ can be chosen as the QSS species in the ROBER problem, so its algebraic equation can be obtained by assuming $\mathrm{d}y_{2}/\mathrm{d}t=0$. Because QSS species are usually radicals and unstable intermediates with short time scales, the elimination of their ODEs can greatly reduce the stiffness. The stiff-PINNs were used to solve the ROBER and POLLU problems with one and ten QSS species, respectively, and yielded satisfactory results. De Florio et al. 262 used X-TFC 136 (see Section 2.3.2) to solve stiff chemical kinetic ODEs, where the TFC can guarantee the ICs are exactly satisfied, as illustrated in Figure 6. The temporal domain was decomposed into several intervals to deal with the long-time issue. The X-TFC showed great performance for 4 tested problems, and even outperformed the traditional numerical methods in some cases. Wang et al. 263 used PINNs to solve the ROBER problem. Hard constraints on ICs via logarithmic functions were employed, yielding accurate results. Galaris et al. 264 proposed the physics-informed random projection NNs (PI-RPNNs) to solve stiff ODEs, where the single-layer NN had fixed weights between the input and hidden layer so only the parameters from the hidden to the output layer needed to be optimized. The loss functions were minimized by the Gauss-Newton method and the hard constraints on ICs were designed. Four stiff ODE problems were tested and PI-RPNN yielded good accuracy and can outperform the traditional solvers in the cases with steep gradients. Later, Fabiani et al. 265 extended the PI-RPNNs to other stiff ODEs and three index-1 DAEs. For more practical applications, Li et al. 266 applied PINNs to solve the coal gasification chemical kinetic problems. The unreacted-core shrinking model was adopted for the gas-solid reactions, the temperature was simplified to be constant, and the ICs were hardcoded. Results showed that PINNs could accurately predict the production rates of the gaseous species.
@@
Figure 5: Schematic of the stiff-PINN for the ROBER problem. Reprinted from Reference 261, with permission.
@@
Figure 6: Schematic of the X-TFC algorithm to solve stiff chemical kinetics equations. Reprinted from Reference 262, with permission.
@@
Although the above studied focus on solving the chemical kinetic ODEs in a forward way without data supervision, inverse problems can also be solved when the available data are utilized. Weng et al. 267 developed the multiscale PINNs for stiff chemical kinetic problems, where the species with different time scales were classified and trained by multiple NNs. By introducing ground truth data points, the proposed method showed superior ability on solving three stiff chemical kinetic problems. Gusmao et al. 268 proposed the kinetics-informed NNs (KINNs) to solve forward and inverse problems of chemical kinetics, as shown in Figure 7. Practical species and reaction types were considered, including three types of species (gas, absorbed molecules, and surface radicals) and five types of reactions (denoted by g, a, d, c, and s). ICs were enforced by an IC operator coupled with a shallow NN and species conservation was enforced by trigonometric transformations. In forward problems, no data were provided, while in inverse problems, synthetic data were available and three scenarios were designed based on the information about chemical intermediates and adsorbed molecules, and the presence of noise. The unknown kinetic parameters were optimized together with the NN parameters in inverse problems. The results showed that KINNs can accurately solve the ODEs in forward problems and estimate the kinetic parameters in inverse problems. Later, KINNs were applied to model the data from temporal analysis of products 286 and extended to a robust version by incorporating the maximum likelihood approach 287. Bibeau et al. 269 used PINNs to predict the kinetics of biodiesel production in microwave reactors. The system contained five species and three reactions. The Arrhenius equation was linearized and the temperature evolution was simplified to a first-order ODE, resulting in 15 unknown parameters related to the reaction kinetics and energy balance, which were estimated by PINNs from experimental data. Results showed that the reaction kinetics of the biodiesel process could be identified successfully by PINNs with satisfactory extrapolation ability.
@@
Figure 7: Kinetics-informed neural networks. Reprinted from Reference 268, with permission.